OpenStudy (anonymous):

Solve using the quadratic formula. Approximate answers to the nearest tenth. x^2-4x-7=0

OpenStudy (mathstudent55):

Do you know the quadratic formula?

OpenStudy (anonymous):

Yeah, but im just confused on how to turn it into a decimal.

OpenStudy (mathstudent55):

What do you have that you need to turn into a decimal?

OpenStudy (anonymous):

well the thing says approximate to the nearest tenth

OpenStudy (mathstudent55):

Yes, I understand. Once you used the quadratic formula, what solutions did you get?

OpenStudy (mathstudent55):

First use the quadratic formula. You'll get two solutions with square roots in them. Then use a calculator to find decimal approxinmations of the solutions.

OpenStudy (anonymous):

Okay I did that, but how do I write it?

OpenStudy (anonymous):

cuz the decimal is 3.32

OpenStudy (anonymous):

\[\left\{x=2-\sqrt{11},x=2+\sqrt{11}\right\} \]\[\{x=-1.3,x=5.3\} \]

OpenStudy (mathstudent55):

What did you get as solutions of the quadratic formula?

OpenStudy (mathstudent55):